OpenStudy (anonymous):

Estimate by rounding to the nearest half. 3 3/7 + 4 7/8 - 2 5/9 choices: 6 5 1/2 4 6 1/2

Parth (parthkohli):

3 1/2 + 5 - 2 1/2 = 1 + 5 = 6

OpenStudy (anonymous):

I'M DONE!!!!!!!!!!! THANK YOU!!